# Towards Inclusive ASR: Investigating Voice Conversion for Dysarthric Speech Recognition in Low-Resource Languages

### v3

@@ 1 Introduction
Dysarthria is a motor speech disorder caused by neurological conditions such as Cerebral Palsy (CP), Parkinson’s disease (PD), and Amyotrophic Lateral Sclerosis (ALS) [1]. These neurological impairments affect the coordination and strength of the muscles involved in speech production, resulting in reduced speech intelligibility [2]. Consequently, despite notable progress, automatic speech recognition (ASR) systems still struggle to process dysarthric speech [3].
@@ 3.1 Datasets
PC-GITA [28] includes 50 patients with Parkinson’s Disease (PD) and 50 healthy subjects evenly matched in age and gender, all native Spanish speakers from Colombia. Each participant read 10 phonetically balanced sentences and 25 isolated words representing the phonetic inventory of Colombian Spanish.
EasyCall [29] is an Italian dysarthric speech dataset comprising 21 healthy speakers and 26 dysarthric speakers. The dysarthric group includes individuals with conditions such as PD, Huntington’s Disease, ALS, peripheral neuropathy, and myopathic or myasthenic lesions. Each speaker recorded 66 to 69 mobile commands, including isolated words and short phrases.
@@ 3.4 Automatic Speech Recognition System
We considered MMS-1B-ALL model [23] for all ASR experiments. The model’s adapter was fine-tuned for each language on FLEURS, including augmented versions when available, for five epochs with a 0.0005 learning rate. We evaluated ASR performance across four fine-tuning setups: (1) off-the-shelf without fine-tuning, (2) fine-tuning with healthy speech from FLEURS, (3) fine-tuning with healthy speech and conventional augmentations (speed or tempo perturbation), and (4) fine-tuning with our proposed VC approach (speaker-only VC or speaker-prosody VC). ASR performance was assessed on dysarthric speech datasets described in Section 2, using character error rate (CER) as an evalaution metric, with separate evaluations for word-level and sentence-level utterances. See our codebase [^1] https://github.com/chinjouli/dysaug-vc for more details.
@@ 5 Analyses of Generated Dysarthric Data
ASR performance improvement in Section 4 can attributed to various reasons beyond augmented data being similar to dysarthric speech. To verify whether the generated speech exhibits dysarthric characteristics, we conduct both objective and subjective analyses on the generated data. Unlike in our ASR studies, where the generated data was used for fine-tuning, this section evaluates it as the test set. We pose the following question to both the model and human evaluators: “Does the generated speech sound like dysarthria?”
@@ 5.1 Objective Evaluation: Classification
Table 3 presents the proportion of audio samples classified as dysarthric relative to the total number of samples, evaluated under different augmentation techniques. Results on FLEURS with no augmentation (“None”) serve as the baseline, since our generated data originates from FLEURS. Speed and tempo perturbations had minimal impact, suggesting these methods do not effectively simulate dysarthria, aligning with our ASR results. In contrast, VC-based augmentations led to a substantial increase in the proportion of samples classified as dysarthric, confirming that speaker and prosodic modifications enhance dysarthric-like characteristics. Speaker-prosody VC generally outperformed speaker-only VC, reinforcing the role of prosodic transfer in dysarthria simulation. Notably, Italian samples, including FLEURS, were frequently classified as dysarthric, suggesting potential noise in the FLEURS Italian dataset, warranting further investigation.
@@ 5.2 Subjective Evaluation: Perceptual Tests
For the subjective evaluations, we selected 10 audio samples from each group: healthy and dysarthric speech from the dysarthric datasets, and generated speech from each augmentation method. The samples were balanced across severity levels and randomly selected within each. For each language, two native speakers with no prior training in speech pathology provided judgments in two tasks. We provided information on perceptual characteristics of dysarthric speech before evaluation1. The first evaluation was conducted on dysarthric datasets, where participants rated whether the audio sounded dysarthric (“Dys.?”) using a 1-to-4 Likert scale [40], ranging from “definitely healthy” to “definitely dysarthric.” As the second evaluation, we asked the participants to apply the same scale to generated speech samples, asking how closely generated speech approximates the dysarthric or healthy speech characteristics. For the generated samples, we asked additional question on the naturalness in the target language (“Your Lang.?”). 1-to-4 Likert scale was used, ranging from “Not at all my language” to “Definitely my language.” This helps to provide information on how well the linguistic characteristics were preserved by the augmentation methods.
Table 4 shows that perceptual tests yield similar trends to objective evaluation. Speaker-prosody VC consistently obtained the highest dysarthria ratings, indicating a strongest shift toward dysarthric traits. Speaker-only VC alone also increased dysarthria perception but to a lesser degree, while speed and tempo-based data augmentations had minimal impact relative to healthy data. Despite its effectiveness in introducing dysarthric characteristics, speaker-prosody VC showed the lowest linguistic similarity ratings, indicating reduced naturalness. By contrast, speed and tempo-based augmentations best maintained linguistic characteristics, with speaker-only VC providing an intermediate balance. These findings underscore a trade-off between enhancing dysarthric characteristics and preserving linguistic naturalness, highlighting the need to balance augmentation strategies according to specific application requirements. For VC methods, multiple evaluators mentioned, “They sound like Americans trying to speak my language,” implying future directions to mitigate such artifacts.

### v4

@@ 1 Introduction
Dysarthria is a motor speech disorder caused by neurological conditions such as Cerebral Palsy (CP), Parkinson's disease (PD), and Amyotrophic Lateral Sclerosis (ALS) [1]. These neurological impairments affect the coordination and strength of the muscles involved in speech production, resulting in reduced speech intelligibility [2]. Consequently, despite notable progress, automatic speech recognition (ASR) systems still struggle to process dysarthric speech [3].
@@ 3.1 Datasets
PC-GITA [28] includes 50 patients with Parkinson's Disease (PD) and 50 healthy subjects evenly matched in age and gender, all native Spanish speakers from Colombia. Each participant read 10 phonetically balanced sentences and 25 isolated words representing the phonetic inventory of Colombian Spanish.
EasyCall [29] is an Italian dysarthric speech dataset comprising 21 healthy speakers and 26 dysarthric speakers. The dysarthric group includes individuals with conditions such as PD, Huntington's Disease, ALS, peripheral neuropathy, and myopathic or myasthenic lesions. Each speaker recorded 66 to 69 mobile commands, including isolated words and short phrases.
@@ 3.4 Automatic Speech Recognition System
We considered MMS-1B-ALL model [23] for all ASR experiments. The model's adapter was fine-tuned for each language on FLEURS, including augmented versions when available, for five epochs with a 0.0005 learning rate. We evaluated ASR performance across four fine-tuning setups: (1) off-the-shelf without fine-tuning, (2) fine-tuning with healthy speech from FLEURS, (3) fine-tuning with healthy speech and conventional augmentations (speed or tempo perturbation), and (4) fine-tuning with our proposed VC approach (speaker-only VC or speaker-prosody VC). ASR performance was assessed on dysarthric speech datasets described in Section 2, using character error rate (CER) as an evalaution metric, with separate evaluations for word-level and sentence-level utterances. See our codebase [^1] https://github.com/chinjouli/dysaug-vc for more details.
@@ 5 Analyses of Generated Dysarthric Data
ASR performance improvement in Section 4 can attributed to various reasons beyond augmented data being similar to dysarthric speech. To verify whether the generated speech exhibits dysarthric characteristics, we conduct both objective and subjective analyses on the generated data. Unlike in our ASR studies, where the generated data was used for fine-tuning, this section evaluates it as the test set. We pose the following question to both the model and human evaluators: ``Does the generated speech sound like dysarthria?''
@@ 5.1 Objective Evaluation: Classification
Table 3 presents the proportion of audio samples classified as dysarthric relative to the total number of samples, evaluated under different augmentation techniques. Results on FLEURS with no augmentation (``None'') serve as the baseline, since our generated data originates from FLEURS. Speed and tempo perturbations had minimal impact, suggesting these methods do not effectively simulate dysarthria, aligning with our ASR results. In contrast, VC-based augmentations led to a substantial increase in the proportion of samples classified as dysarthric, confirming that speaker and prosodic modifications enhance dysarthric-like characteristics. Speaker-prosody VC generally outperformed speaker-only VC, reinforcing the role of prosodic transfer in dysarthria simulation. Notably, Italian samples, including FLEURS, were frequently classified as dysarthric, suggesting potential noise in the FLEURS Italian dataset, warranting further investigation.
@@ 5.2 Subjective Evaluation: Perceptual Tests
For the subjective evaluations, we selected 10 audio samples from each group: healthy and dysarthric speech from the dysarthric datasets, and generated speech from each augmentation method. The samples were balanced across severity levels and randomly selected within each. For each language, two native speakers with no prior training in speech pathology provided judgments in two tasks. We provided information on perceptual characteristics of dysarthric speech before evaluation1. The first evaluation was conducted on dysarthric datasets, where participants rated whether the audio sounded dysarthric (``Dys.?'') using a 1-to-4 Likert scale [40], ranging from ``definitely healthy'' to ``definitely dysarthric.'' As the second evaluation, we asked the participants to apply the same scale to generated speech samples, asking how closely generated speech approximates the dysarthric or healthy speech characteristics. For the generated samples, we asked additional question on the naturalness in the target language (``Your Lang.?''). 1-to-4 Likert scale was used, ranging from ``Not at all my language'' to ``Definitely my language.'' This helps to provide information on how well the linguistic characteristics were preserved by the augmentation methods.
Table 4 shows that perceptual tests yield similar trends to objective evaluation. Speaker-prosody VC consistently obtained the highest dysarthria ratings, indicating a strongest shift toward dysarthric traits. Speaker-only VC alone also increased dysarthria perception but to a lesser degree, while speed and tempo-based data augmentations had minimal impact relative to healthy data. Despite its effectiveness in introducing dysarthric characteristics, speaker-prosody VC showed the lowest linguistic similarity ratings, indicating reduced naturalness. By contrast, speed and tempo-based augmentations best maintained linguistic characteristics, with speaker-only VC providing an intermediate balance. These findings underscore a trade-off between enhancing dysarthric characteristics and preserving linguistic naturalness, highlighting the need to balance augmentation strategies according to specific application requirements. For VC methods, multiple evaluators mentioned, ``They sound like Americans trying to speak my language,'' implying future directions to mitigate such artifacts.